OpenStudy (anonymous):

Someone please help. solve and check 2/(1-x) = 3/(1+x)

OpenStudy (anonymous):

x=1/5

OpenStudy (anonymous):

This is what I got. Am I correct? 2/(1-x) = 3/(1+x) 2(1+x) = 3(1-x) 2 + 2x = 3 - 3x 2x + 3x = 3 - 2 5x = 1 x = 1/5

OpenStudy (anonymous):

x=0.2

OpenStudy (anonymous):

your works correct